OpenStudy (anonymous):

Write the following in standard form using integers. y = 2/3x - 7

OpenStudy (caitlinnr14):

the std form is ax + by = c y = 2/3x -7 multiply both sides by 3 3y = 2x - 21 2x - 3y = 21

OpenStudy (anonymous):

Why do you multiply both sides by 3?